A company will need $75,000 in 5 years. To meet this goal the company deposits money in an account today that pays 11% annual interest compounded quarterly. What amount should be invested to total $75,000 in 5 years?
1
Expert's answer
2020-11-09T20:31:21-0500

"\bold {Answer}"

Amount to be invested today to total "\$75, 000" in 5 years:


"PV = \$43,593.79"



"\bold {Solution}"

We are provided:

Future value, "FV = \$75,000"

Annual interest rate, "i\% = \$11\%"

Number of compounding periods, "m = 4"

Number of years, "n = 5 \space years"


Required:

Present value, "PV = ?"


The formular for future value is given by: "FV = PV \left (1 + \dfrac {i}{m}\right)^{mn}"


"=> \$75,000 = PV \left (1 + \dfrac {0.11}{4}\right)^{(4×5)}"



"=> \$75,000 = PV \left (1 + 0.0275 \right)^{20}"


"=> \$75,000 = PV \left (1.0275 \right)^{20}"


"=> PV =\dfrac { \$75,000}{1.0275^{20}}"


"= \dfrac {\$75,000}{1.7204284313}"


"= \$43,593.792474"

"=\bold {\$43,593.79}"